OpenStudy (anonymous):

need help... Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8)

OpenStudy (phi):

can you find the slope between the two points ?

OpenStudy (anonymous):

yes....-8+3/4-7=-5/-3

OpenStudy (phi):

and -5/-3 is 5/3 (the two minus signs make a plus) now use the point slope formula y - y0 = m ( x - x0) where (x0,y0) is a point you know that is on the line, and m is the slope = 5/3 for this problem can you replace x0, y0 and m with the correct numbers? what do you get ?

OpenStudy (anonymous):

y+3=5/3(x-7)

OpenStudy (phi):

yes... of course you could have used the other point ... in case you were wondering. now we need to make it into standard form. there are no fractions in standard form. to get rid of the 5/3, multiply both sides by 3, like this: \[ 3(y+3) = 3 \cdot \frac{5}{3} \cdot (x-7)\] can you distribute the 3 on the left side and the 5 on the right side ?

OpenStudy (phi):

in case you did not know... 3/3 "cancels" (becomes 1, and multiply by 1 can be ignored)

OpenStudy (phi):

distribute means multiply each term inside the parens. for example 3(y+3) means 3*y + 3*3 or 3y + 9

OpenStudy (anonymous):

ok....3y+9=5x-35

OpenStudy (phi):

any ideas what do to next ?

OpenStudy (anonymous):

+ 5x on both sides??????????

OpenStudy (phi):

almost. you want the 5x on the right side to "go away" you do that by adding -5x because 5x take away 5x leaves nothing... of course if you add -5x to the right side, you have to add -5x to the left side also. what do you get?

OpenStudy (anonymous):

will it be -5x+3y+9=-35

OpenStudy (phi):

yes

OpenStudy (phi):

3y+9=5x-35 -5x+3y +9 = 5x - 5x -35 or -5x + 3y + 9 = -35 standard form is Ax + By = C (and A is positive) you are almost done. you need to "move" the 9 to the other side. To do that add -9 to both sides, and simplify

OpenStudy (anonymous):

so it will be 5x+3y=-26

OpenStudy (phi):

one step at a time -5x + 3y + 9 = -35 add -9 to both sides

OpenStudy (anonymous):

-5x+3y+-9=-35+-9 -5x+3y=-44?????????

OpenStudy (phi):

yes, but I would write it like this first, the original equation: -5x + 3y + 9 = -35 then write in -9 on both sides: -5x + 3y + 9 + -9 = -35+-9 then simplify: 9-9 is 0 on the left side: -5x + 3y = -35 + -9 and -35 - 9 is -44 on the right side: -5x+3y=-44

OpenStudy (phi):

now the last step. we have -5x+3y=-44 which matches Ax + By = C but the A is supposed to be positive. We have -5 which is negative multiply both sides by -1: -1(-5x+3y) = -1* -44 distribute the -1 on the left what do you get ?

OpenStudy (anonymous):

5x+3y=44

OpenStudy (phi):

almost. distribute the -1 means multiply each term inside the parens by -1 try again -1(-5x+3y) = 44

OpenStudy (anonymous):

-1*-5=5 -1*3=-3 than-44*-1=44 which is 5x+-3y=44????

OpenStudy (phi):

yes, exactly. but generally people write 5x + -3y as 5x - 3y (it means the same thing, but it looks nicer) 5x - 3y = 44

OpenStudy (anonymous):

ok thank you for all the help.......:)

OpenStudy (phi):

just to make sure we didn't make a mistake, we can test the other point (4, -8) in this equation: 5x - 3y = 44 5*4 + -3*-8 =? 44 20 +24 =? 44 44 = ? 44 yes, it worked.

OpenStudy (anonymous):

ok....